OpenStudy (anonymous):

At our school, 5 out of 7 students is involved in an after school activity. How many of the 700 students are in an after school activity at our school?

OpenStudy (anonymous):

\[700 \cdot \frac{5}{7} = 5 \cdot 100 = 500\]

OpenStudy (anonymous):

Thank you.